OpenStudy (anonymous):

HELP! PLEASE! QUADRATIC EQUATIONS! Solve the equation by finding the square root. 3a^2 + 12 = 0

OpenStudy (anonymous):

3a^2=-12

OpenStudy (anonymous):

divide both sides by 3

OpenStudy (anonymous):

good luck solving this one

OpenStudy (31356):

Yes, lol :D

OpenStudy (anonymous):

\(3a^2\geq 0\) so \(3a^2+12\geq 12\) so you can't solve \(3a^2+12=0\)

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

Thanks for the help though guys! @esamalaa @satellite73 @31356

OpenStudy (31356):

My pleasure, although I didn't help much... lol :D